# ToolACE: Winning the Points of LLM Function Calling

### v1

@@ 5.3.2 Ablation on Complexity
Data Sampling for Various Complexity. To effectively assess the impact of dataset complexity on the model’s performance, we have conducted a sampling of the entire dataset based on the aforementioned complexity assessment metrics. This process has yielded three distinct subsets of varying complexity levels: $\text{ToolACE}_{easy}$, $\text{ToolACE}_{medium}$, and $\text{ToolACE}_{hard}$, each containing about 60,000 instances. Figure 18 and Figure 19 in Appendix C illustrate the different distributions of these subsets in terms of Dataset-level Complexity and Query-level Complexity. The rationale behind this stratified sampling approach is to create a controlled environment where the influence of complexity can be systematically analyzed. By maintaining equal sample sizes across subsets, we ensure a fair comparison while varying the complexity, which allows for a more nuanced understanding of how complexity affects model performance.
Effects of Complexity. We conduct experiments on those three subsets with varying complexity and evaluate the fine-tuned models on the BFCL benchmark. The results are illustrated in Figure 13. The model trained on $\text{ToolACE}_{medium}$ shows slight superiority compared with another two subsets, for both overall and tool-use accuracy. This finding aligns with our hypothesis that optimal data complexity is essential for LLM training; excessively simple or complex data can hinder performance.
@@
(c) Relevance Accuracy
(d) Overall Accuracy
@@ 5.3.3 Ablation on Diversity
Data Sampling for Various Diversity. To assess the impacts of the diversity, we employed a sampling strategy to generate three subsets with varying degrees of diversity, namely $\text{ToolACE}_{low}$, $\text{ToolACE}_{medium}$, and $\text{ToolACE}_{high}$. Initially, all APIs are clustered into 30 groups using K-means based on their names and descriptions. Subsequently, API sets are constructed by selecting APIs from 6, 14, and 30 clusters, respectively. Instances are then categorized into three subsets according to their associated APIs. Approximately 30,000 instances are randomly selected from each subset, resulting in three training sets with distinct levels of diversity.
Effects of Diversity. Experiments are conducted to train LLaMA-3.1-8B-Instruct on three subsets described above. The results on BFCL are reported in Figure 13. A clear correlation between training data diversity and overall model accuracy is observed, emphasizing the critical role of API diversity in model performance. Notably, improvements in relevance detection are particularly pronounced, suggesting that exposure to a wider range of APIs enhances the model’s ability to discriminate between subtle API differences, thereby enhancing the ability of irrelevance detection.
@@ 5.4 Scaling Performance of Model Size
Scaling laws posit a correlation between model size and performance. To investigate the scalability of functional calling performance, we conduct experiments using the Qwen-1.5-xB-Chat series, which offers a range of model sizes (0.5B, 1.8B, 4B, 7B, etc.). Both raw and fine-tuned (using our dataset) models are evaluated on the BFCL benchmark, with results presented in Figure 14. As expected, larger models exhibit superior performance in functional calling tasks, as evidenced by improvements in AST accuracy and executable accuracy. Smaller raw models demonstrate unexpectedly high relevance detection scores, likely attributable to the generation of unparsable outputs due to limited instruction-following capabilities. Conversely, fine-tuned models display consistent scaling behavior across all evaluation metrics, highlighting the potential of the ToolACE to enhance the performance of larger LLMs.
@@ 5.5 Study on Various Backbone LLMs
To investigate the influence of the LLM backbone, we experiment with several (approximately) 8B-scale models: Qwen1.5-7B-Chat [3], LLaMA-3-8B-Instruct, and LLaMA-3.1-8B-Instruct. Fine-tuned models are evaluated on the BFCL benchmark, with results presented in Figure 16. Across all models, fine-tuning yields substantial performance gains, highlighting the effectiveness of our ToolACE. Due to differences in pre-training corpora, such as Qwen is trained with more Chinese conversational samples, raw models exhibit varying functional calling capabilities, with LLaMA-3.1-8B-Instruct demonstrating superior performance. While this hierarchy persisted after fine-tuning, the performance gaps narrowed, suggesting that our dataset can potentially enhance the functional-calling abilities of those LLMs tailored for other skills, such as conversational skills.
@@
Figure 16: Comparisons of general capabilities.
@@ 5.6 Study on General Capabilities
To assess the impact of ToolACE training on broader capabilities of LLMs, we conduct experiments across multiple benchmarks evaluating general ability (MMLU [7, 8]), coding (HumanEval [4]), mathematics (GSM8K [5]), reasoning (CommonSenseQA [18]), and functional calling (BFCL [23]). Raw LLaMA-3-8B-Instruct, LLaMA-3.1-8B-Instruct, and the functionally specialized xLAM-7B-fc-r serve as baselines. Results are presented in Figure 16. ToolACE-8B exhibits substantial improvements over xLAM-7B-fc-r across most benchmarks, with particularly pronounced gains in MMLU, GSM8K, and CommonSenseQA. Compared to the raw LLaMA-3.1-8B-Instruct, ToolACE-8B demonstrates negligible performance degradation on some benchmarks while achieving significant enhancements in functional calling. These findings suggest that the ToolACE dataset effectively enhances functional calling capabilities without compromising the underlying LLM’s general abilities.
@@ 6 Conclusion
This paper presents ToolACE, an automated data generation pipeline designed to enhance the function-calling capabilities of large language models. ToolACE utilizes a novel self-evolution synthesis process and a multi-agent interactive system to curate accurate, complex, and diverse APIs and dialogs. Our results demonstrate that even smaller models trained with ToolACE can achieve state-of-the-art performance, thereby advancing the field and setting new benchmarks for tool-augmented AI agents.
7 Acknowledgements
We thank the contribution of Mei Li, Wei Tang, Xi Wang, Meide Zhang, Ke Xiong, and Rencong Shi to this work.
Appendix A Rule Examples in Rule Verification Layer
Table 6 lists the check rules we have applied. For instance, to verify function calling executability, we implement the following procedures: First, we confirm that the API name matches one from the given tool list. Next, we verify that all required parameters are accurately provided. Finally, we use regular expressions to ensure that the parameter formats and patterns adhere to those specified in the API documentation. These procedures allow us to validate the correctness and executability of function calls without actual execution, which enhances efficiency and reduces deployment overhead.
@@ Appendix A Rule Examples in Rule Verification Layer
Function Calling Executability Check if the API name is in the tool list.
@@
Data Sample Consistency Check if the API names in the function call and the tool response are consistent.
@@
Appendix B An example of Different API formats
Figure 17 displays an example of API definitions with different formats, including original Json, XML, Markdown and natural language.
Figure 17: Different format of API definitions.
Appendix C Data Distribution on Complexity Ablation Study
Figure 18 and Figure 19 show the distribution of the three sampled subsets concerning complexity.
Figure 18: Data distributions of subsets with different complexity.
Figure 19: Data distribution of subset scores.
@@ Appendix D Case Study
Figure 20 presents a data sample of parallel function calling. This type of data focuses on building the model’s ability to understanding and extract key parameters from the user query, which makes models learn to call the correct function repeatly for accomplishing a task. In this example, the query indicates that the user needs to get the event information of Theatre, Dance, and Music during 2021-04-01 and 2021-05-01 respectively. The assistant correctly recognizes that it needs to call performanceArt.get_upcoming_events for three times with different assignments of the parameter "category".

### v2

@@ 3.3.2 Ablation on Complexity
Data Sampling for Various Complexity. To effectively assess the impact of dataset complexity on the model’s performance, we have conducted a sampling of the entire dataset based on the aforementioned complexity assessment metrics. We compute and sort the complexity for each data sample using Eq. (1), and select the bottom, middle, and top 60,000 instancess as $\text{ToolACE}_{easy}$, $\text{ToolACE}_{medium}$, $\text{ToolACE}_{hard}$, respectively, yielding three distinct subsets of varying complexity levels
The rationale behind this stratified sampling approach is to create a controlled environment where the influence of complexity can be systematically analyzed. By maintaining equal sample sizes across subsets, we ensure a fair comparison while varying the complexity, which allows for a more nuanced understanding of how complexity affects model performance.
Effects of Complexity. We conduct experiments by training LLaMA-3.1-8B-Instruct with those three subsets with varying complexity and evaluate the fine-tuned models on the BFCL benchmark. The results are illustrated in Figure 5. The model trained on $\text{ToolACE}_{medium}$ shows slight superiority compared with another two subsets, for both overall and tool-use accuracy. This finding aligns with our hypothesis that optimal data complexity is essential for LLM training, as data that is either too simple or overly complex can prevent the model from reaching its full performance potential.
@@
(c) Overall Accuracy
@@ 3.3.3 Ablation on Diversity
Data Sampling for Various Diversity. To assess the impacts of the diversity, we generate three subsets with varying diversity, namely $\text{ToolACE}_{low}$, $\text{ToolACE}_{medium}$, and $\text{ToolACE}_{high}$. Initially, all APIs are clustered into 30 groups based on the API context tree. Subsequently, three API subsets are constructed by selecting APIs from 6, 14, and 30 clusters, respectively. Instances are then categorized into three subsets according to their associated APIs. Approximately 30,000 instances are randomly selected from each subset, resulting in three training sets with distinct levels of diversity.
Effects of Diversity. Experiments are conducted to train LLaMA-3.1-8B-Instruct on three subsets described above. The results on BFCL are reported in Figure 5. A positive correlation between training data diversity and overall model accuracy is observed, emphasizing the critical role of API diversity in model performance. Notably, improvements in relevance detection are particularly pronounced, suggesting that exposure to a wider range of APIs enhances the model’s ability to discriminate between subtle API differences, thereby enhancing the ability of irrelevance detection.
@@ 3.4 Scaling Performance of Model Size
Scaling laws posit a correlation between model size and performance. To investigate the scalability of functional calling capabilities, we conduct experiments using the Qwen-1.5-xB-Chat series, which includes a range of model sizes (0.5B, 1.8B, 4B, 7B, etc.). Both raw and fine-tuned (using our dataset) models are evaluated on the BFCL benchmark, with results presented in Figure 6. As expected, larger models exhibit superior performance in functional calling, as evidenced by improvements in both AST and Executable accuracy. While smaller raw models (0.5B and 1.8B) showed minimal function-calling ability, struggling to generate structured outputs, fine-tuning on the ToolACE dataset significantly enhanced their capabilities. The fine-tuned models exhibit consistent scaling performance, highlighting the potential of ToolACE to boost the performance of larger LLMs.
@@ 3.5 Study on Various Backbone LLMs
To investigate the influence of the LLM backbone, we experiment with several (approximately) 8B-scale models: Qwen1.5-7B-Chat Bai et al. (2023), LLaMA-3-8B-Instruct, and LLaMA-3.1-8B-Instruct. Fine-tuned models are evaluated on the BFCL benchmark, with results presented in Figure 8. Across all models, fine-tuning yields substantial performance gains, highlighting the effectiveness of our ToolACE. Due to differences in pre-training corpora, such as Qwen is trained with more Chinese conversational samples, raw models exhibit varying functional calling capabilities, with LLaMA-3.1-8B-Instruct demonstrating superior performance. While this hierarchy persisted after fine-tuning, the performance gaps narrowed, suggesting that our dataset can potentially enhance the functional-calling abilities of those LLMs tailored for other skills, such as conversational skills.
@@
Figure 8: General capabilities.
@@ 3.6 Study on General Capabilities
To assess the impact of ToolACE training on broader capabilities of LLMs, we conduct experiments across multiple benchmarks evaluating general ability, including MMLU Hendrycks et al. (2021a; b), HumanEval Chen et al. (2021) (coding), GSM8K Cobbe et al. (2021) (mathematics), CommonSenseQA Talmor et al. (2019) (reasoning), and BFCL Yan et al. (2024) (functional calling). Raw LLaMA-3-8B-Instruct, LLaMA-3.1-8B-Instruct, functionally specialized xLAM-7B-fc-r, and GPT-4 serve as baselines. Results are presented in Figure 8. ToolACE-8B substantially improves over xLAM-7B-fc-r across most benchmarks, with particularly pronounced gains in MMLU, GSM8K, and CommonSenseQA. Compared to GPT-4, ToolACE-8B shows clear limitations in reasoning and understanding. This is primarily due to the scale of the model and its training corpus. Compared to the raw LLaMA-3.1-8B-Instruct, ToolACE-8B demonstrates negligible performance degradation on some benchmarks while achieving significant enhancements in functional calling. These findings suggest that the ToolACE dataset effectively enhances functional calling capabilities without compromising the underlying LLM’s general abilities. This success highlights the potential of specialized models in one specific domain, the challenge of simultaneously enhancing multiple capabilities, alongside functional-calling performance, remains an open question. The detailed analysis of the limitations can be referred to in Appendix H.
4 Related Work
Tool Learning. Integrating external tools allows LLMs to expand the boundaries of their capabilities, enabling more specialized, precise, and dependable problem-solving (Qin et al., 2023). Methods for equipping LLMs with tool-use capabilities generally fall into two types: tuning-free approaches and tool-augmented tuning. Tuning-free methods let LLMs use tools by providing in-context tool descriptions and examples, requiring no additional training Mialon et al. (2023); Hsieh et al. (2023); Ruan et al. (2023). A well-known technique is ReAct Yao et al. (2023), which enables LLMs to alternate between reasoning and actions to solve complex tasks. However, as these approaches depend heavily on the model’s initial abilities, tool-augmented tuning has gained more attention for directly improving tool use Qin et al. (2023); Schick et al. (2023); Patil et al. (2023); Tang et al. (2023); Liu et al. (2024); Abdelaziz et al. (2024). Many of these methods rely on existing APIs but lack robust systems for generating and validating data. Our ToolACE overcomes this limitation by implementing a well-designed pipeline that ensures greater diversity, complexity, and accuracy.
Data Synthesis. As LLMs grow more advanced, relying solely on existing human-generated data becomes insufficient for further progress Bauer et al. (2024). A key strategy involves modifying or augmenting datasets using specialized prompting techniques Wang et al. (2023); Xu et al. (2023); Yu et al. (2023). Given the scarcity of tool-use datasets, Basu et al. (2024) repurpose data from other domains for tool-use applications, while others Qin et al. (2023); Tang et al. (2023); Liu et al. (2024) depend on publicly available APIs, often producing single-turn instructions with basic tool interactions. ToolACE offers a more comprehensive approach, incorporating both tool synthesis and dialogue generation, along with a verification module to ensure data quality.
@@ 5 Conclusion
This paper presents ToolACE, an automated data generation pipeline developed to enhance the function-calling capabilities of large language models. ToolACE employs a novel self-evolution synthesis process and a self-guided data generation method to curate accurate, complex, and diverse synthetic APIs and dialogs. Our results demonstrate that even smaller models trained with ToolACE can achieve state-of-the-art performance, thereby advancing the field and setting new benchmarks for tool-augmented AI agents.
Appendix A An example subtree of the API context tree for the Entertainment domain.
Figure 9: A subtree of the constructed API context tree for the Entertainment domain.
Appendix B Rule Examples in Rule Verification Layer
Table 4 outlines the check rules we use, which consists of four aspects: API definition clarity, function calling executability, dialog correctness, and data sample consistency.
@@ Appendix B Rule Examples in Rule Verification Layer
Function Calling
Executability
Check if the API name is in the tool list.
@@
Data Sample
Consistency
Check if the API names in the function call and the tool response are consistent.
@@
Appendix C Experimental Details
C.1 Benchmarks
BFCL.
Berkeley Function-Calling Benchmark (BFCL) Yan et al. (2024) is a comprehensive evaluation framework for assessing the function-calling capabilities of LLMs across various languages, application domains, and complex use cases. BFCL covers tasks including multiple function calls, parallel function calls, multi-turn function calls, and multi-step function calls. BFCL contains 4,951 test cases: 3,951 single-turn cases and 1,000 multi-turn cases, focusing on dynamic, real-world scenarios.
BFCL evaluates multiple function calling tasks using the following metrics:
  • Abstract Syntax Tree (AST) Evaluation: AST evaluation compares the abstract syntax tree of the function output to the ground truth and the function definition. It captures the correctness of matching the functions, required parameters, parameter types, and values.
  • Executable Function Evaluation: Executable function evaluation assesses the accuracy of the generated API call by executing it and comparing the output with the ground-truth output.
  • Irrelevance: Irrelevance measures the model’s ability to refrain from making function calls given irrelevant user queries. The irrelevance score is calculated as the number of correct non-function-call predictions divided by the total number of test samples.
  • Relevance: Relevance evaluates the model’s ability to output function calls relevant to the user query. In this category, the correctness of the parameter values is not considered. The relevance score is calculated as the number of correct function-call predictions divided by the total number of test samples.
  • Overall Accuracy: Overall accuracy is the unweighted average of the accuracies across all sub-categories.
API-Bank.
API-Bank Li et al. (2023) consists of 314 tool-use dialogues with 753 API calls to assess LLMs’ capabilities in planning, retrieving, and calling APIs, with 363 single calls and 122 multiple calls. API-Bank assesses LLM performance across three capabilities:
  • Call: The ability to call an API based on a given query when the APIs are known.
  • Retrieval+Call: The ability to retrieve and call a single API when the APIs are unknown.
  • Plan+Retrieval+Call: The ability to continuously plan, retrieve, and call multiple APIs when the APIs are unknown.
The evaluation metric for API-Bank is accuracy, calculated as the number of correct predictions divided by the total number of predictions.
C.2 Hyper-parameters
The hyper-parameters of the training process are illustrated in Table 5.
Table 5: Hyper-parameters in experiments for training.
Learning
Rate
WarmUp
Ratio
LR
Scheduler
Batch
Size
Epochs
LoRA
rank
LoRA
alpha
$10^{-4}$ 0.10.10.1 cosine 48 3 16 32
@@ Appendix D Case Study
Figure 10 presents a data sample of parallel function calling. This type of data focuses on building the model’s ability to understand and extract key parameters from the user query, which makes models learn to call the correct function repeatedly for accomplishing a task. In this example, the query indicates that the user needs to get the event information of Theatre, Dance, and Music during 2021-04-01 and 2021-05-01 respectively. The assistant correctly recognizes that it needs to call performanceArt.get_upcoming_events for three times with different assignments of the parameter "category".
@@
Appendix E Additional Experiments
E.1 Comparison with other training data
To further highlight the value of our ToolACE dataset, we conducted additional experiments on two representative datasets (ToolLLM and xLAM), as summarized in Table 6. Specifically, we trained models using the amount of data (25,000 samples) and the same base model (LLaMA-3.1-8B-Instruct) to ensure a fair comparison. The trained models were then evaluated on the BFCL benchmark. The results show that the model trained with our dataset consistently outperforms the others across all categories, further validating the effectiveness of our approach. Notably, the model trained on the xLAM dataset exhibits relatively poor performance in irrelevance detection, likely due to a lack of diverse sample types, such as cases where provided tools cannot solve the task. Moreover, the ToolLLM dataset, which primarily focuses on multi-step and dependent cases, demonstrates weak generalization on the BFCL benchmark.
Table 6: Performances of training with different training datasets. The models are evaluated on the BFCL benchmark.
Training data Overall Non-live(A) Non-live(E) Live(A) Multi turn Rel Irrel
ToolLLM(2.5w) 24.90 42.46 36.36 39.45 0.00 100.00 4.41
xLAM(2.5w) 40.51 81.94 81.77 43.18 4.38 73.17 11.87
ToolACE(2.5w) (Ours) 58.19 86.96 84.73 71.35 16.50 75.61 86.42
E.2 Ablation on various types of data
To underscore the importance of incorporating diverse data types—such as Nested, Parallel, Dependent, and Multi-type, as described in Table 1—we maintain the same overall dataset size (25,000) and selectively replace samples from the Nested, Parallel, Dependent, and Multi-type categories with samples from other data types. Then we train the LLaMA-3.1-8B-Instruct model and evaluate its performance on the BFCL benchmark. The results are summarized in Table 7.
The findings show that removing parallel execution data significantly impairs the model’s ability to invoke multiple tools concurrently. This leads to a notable decrease in performance on Non-live AST and execution tasks, which rely heavily on parallel tool usage. Furthermore, excluding multi-type samples hampers the model’s ability to detect when the candidate tools are irrelevant to the question, resulting in only 6.99% accuracy in irrelevance detection. The model’s ability to handle multi-turn function calls is also impaired. In multi-turn testing, the models sometimes are required not to call functions, but to ask clarifying questions instead.
In contrast, removing nested and dependent samples has a relatively minor effect on the model’s tool-using ability in the BFCL task. Few test samples require nested arguments, and almost none involve dependent tool usage. However, including Dependent and Nested data types contributes to greater data diversity, leading to slight improvements in overall performance.
Table 7: Ablation study on various types of data in ToolACE datasets. The models are evaluated on BFCL benchmark.
Subset Overall Non-live(A) Non-live(E) Live(A) Multi turn Rel Irrel
w.o. Parallel 50.60 74.75 77.30 72.19 1.75 78.05 85.05
w.o. Dependent 57.97 87.63 85.55 71.17 15.50 80.49 85.62
w.o. Nested 57.19 85.46 84.48 70.19 15.38 78.05 86.45
w.o. Multi-type 42.71 89.46 85.50 47.89 1.75 95.12 6.99
ToolACE 58.19 86.96 84.73 71.35 16.50 75.61 86.42
E.3 Ablation on complexity evaluator
To assess the complexity of the training data, we propose a self-guided evaluation method, where the model being trained serves as its own evaluator. To verify the suitability of this approach, we conduct an additional experiment using an independent model (Qwen1.5-7B-Chat, selected for its comparable size to ensure fairness) as the evaluator. The results, shown in Table  8, indicate that using the model being trained as the complexity evaluator offers more accurate guidance, leading to improved performance on the BFCL benchmark. Notably, when the complexity score is assessed using a more advanced model (Qwen-14B), some simpler training samples—those deemed easy by the evaluator but not necessarily by the learner—may be excluded. This leads to slight performance gains on more challenging tasks (e.g., Live AST) but results in degradations on Non-live AST tasks  [^4] Live AST tasks involve rarer and more complex functions compared to Non-live AST tasks, as detailed in BFCL’s documentation.. Conversely, when the evaluator is less capable than the learner, the retained samples tend to be relatively easier for the learner, resulting in improved performance on Non-live AST tasks but a decline in performance on Live AST tasks.
Table 8: Ablation study on complexity evaluator. The evaluator represents the model used to evaluate the complexity. The learner denotes the model to be trained. Qwen-7B, Qwen-14B, and LLaMA-8B are abbreviations of Qwen1.5-7B-Chat, Qwen1.5-14B-Chat, and LLaMA-3.1-8B, respectively.
Evaluator Learner Overall Non-live(A) Non-live(E) Live(A) Multi turn Rel Irrel
Qwen-7B LLaMA-8B 57.61 90.42 85.88 71.30 13.12 87.80 78.12
Qwen-14B LLaMA-8B 57.67 87.98 87.02 73.30 11.75 87.80 84.00
LLaMA-8B LLaMA-8B 59.22 89.27 90.07 73.21 14.37 85.37 83.81
Table 9: Comparison between in-context learning and finetuning.
Method Non-live(A) Non-live(E) Live(A) Rel Irrel
LLaMA-8B (3-shot) 58.81 53.32 36.83 82.93 23.66
ToolACE (finetuning) 89.27 90.07 73.21 85.37 83.81
Appendix F Prompting Templates
To provide a better comprehension of the two benchmarks used in experiments, we have illustrated two examples for BFCL and API-Bank in Figure 13 and Figure 14, respectively.
Appendix G Finetuning VS in-context learning
Given 3 shots for LLaMA-3.1-8B-Instruct, the model still fails to generate correct arguments for such a simple example, such as Figure 16, demonstrating the limited ability in tool using under the in-context learning setting. Besides, due to the addition of few-shot examples, the input to the model consumes a lot more tokens than the fine-tuned model, which successfully addresses the aforementioned example in a zero-shot setting, as presented in Figure 15.
Furthermore, we conducted experiments on BFCL under the RAG-based few-shot in-context learning setting. Specifically, we use the training samples as few-shot examples and retrieve the top 3 most relevant ones according to the user’s question and the provided tools with the BGE model to guide in-context learning. The results illustrated in Table 9 show that few-shot in-context learning not only underperforms fine-tuning in BFCL but also falls short of the zero-shot setting. In many cases, illustrated in Figure 16, the model is misled by the tools in the few-shot examples due to its limited reasoning ability and generalization, selecting those instead of the tools in the test sample, which further exacerbates the model’s hallucination phenomenon.
Appendix H Limitations
While we have conducted extensive experiments to demonstrate the effectiveness of our synthesis dataset in enhancing the functional-calling performance, several challenges remain in our research.
  • Data Complexity Evaluation. The computational complexity of data complexity evaluation is influenced by the size of the model being trained, which limits scalability as both the model size and the number of training samples increase. Additionally, the non-uniform sampling may introduce bias, such as causing the model to struggle with learning difficult examples after one round of training, effectively remaining in its comfort zone. In future work, we will further explore the proposed complexity-based sampling strategy to perform iterative training and sampling over multiple rounds, thereby progressively enhancing the model’s generalization capability on more challenging samples.
  • General Capabilities. Although ToolACE demonstrates comparable performance in functional calling, it still lags behind GPT-4 in other capabilities. While this success highlights the potential of specialized models in one specific domain, the challenge of simultaneously enhancing multiple capabilities, alongside functional-calling performance, remains an open question. Exploring the collaboration of multiple small, domain-specific agents may provide a promising direction for addressing this limitation.
System: You are an expert in composing functions. You are given a question and a set of possible functions. Based on the question, you will need to make one or more function/tool calls to achieve the purpose. If none of the functions can be used, point it out. If the given question lacks the parameters required by the function, also point it out. You should only return the function call in the tools call sections. If you decide to invoke any of the function(s), you MUST put it in the format of [func_name1(params_name1=params_value1, params_name2=params_value2...), func_name2(params)] You SHOULD NOT include any other text in the response. Here is a list of functions in JSON format that you can invoke:
[{"name": "get_weather_data", "description": "Fetches weather data from the Open-Meteo API for the given latitude and longitude.", "parameters": {"type": "dict", "properties": {"coordinates": {"type": "array", "items": {"type": "float"}, "description": "The latitude and longitude of the location."}}, "required": ["coordinates"]}}, {"name": "calc_binomial_probability", "description": "Calculates the probability of getting k successes in n trials.", "parameters": {"type": "dict", "properties": {"n": {"type": "integer", "description": "The number of trials."}, "k": {"type": "float", "description": "The number of successes."}, "p": {"type": "float", "description": "The probability of success."}}, "required": ["n", "k", "p"]}}]
User: I’m planning a small outdoor event in Ottawa, and I need to make sure the weather is going to cooperate. Could you fetch the current weather for me at latitude 45.4215 and longitude -75.6972 using the Open-Meteo API? Also, I’m running a small game at the event, and I’m curious about the chances of winning. If I have 10 attempts at this game and the chance of winning each time is 50%, how likely is it that I’ll win 5 times?
Assistant: [get_weather_data(coordinates=[45.4215, -75.6972]), calc_binomial_probability(n=10, k=5.0, p=0.5)]
Figure 13: Example of BFCL benchmark. The blue text represents the part to be evaluated.
System: Based on the given API description and the existing conversation history 1..t, please generate the API request that the AI should call in step t+1 and output it in the format of [ApiName(key1=’value1’, key2=’value2’, ...)], replace the ApiName with the actual API name, and replace the key and value with the actual parameters. Your output should start with a square bracket "[" and end with a square bracket "]". Do not output any other explanation or prompt or the result of the API call in your output. This year is 2023. Input: User: [User’s plain text or response of API call] AI: [AI’s plain text] ... User: [User’s plain text or response of API call] Expected output: [ApiName(key1=’value1’, key2=’value2’, ...)] API descriptions: [{"name": "GetUserToken", "description": "Get the user token by username and password.", "input_parameters": {"username": {"type": "str", "description": "The username of the user."}, "password": {"type": "str", "description": "The password of the user."}}, "output_parameters": {"token": "type": "str", "description": "The token of the user."}}, {"name": "AddAlarm", "description": "The API for setting an alarm includes a parameter for the alarm time.", "input_parameters": {"token": {"type": "str", "description": "User"s token."}, "time": {"type": "str", "description": "The time for alarm. Format: %Y-%m-%d %H:%M:%S"}}, "output_parameters": {"status": {"type": "str", "description": "success or failed"}}} ]
User: Can you help me set an alarm for 8 am tomorrow?
Assistant: Sure, to set an alarm, I need to authorize your account. Can you please provide your username and password?
User: Sure, my username is foo and password is bar.
Assistant: [GetUserToken(username="foo", passward="bar")]
Tool: [GetUserToken Response: {"token": "z9x8c7v6b5n4m3q2w1"}]
Assistant: Okay, I got your token. What time do you want the alarm to be set for?
User: 8 am tomorrow. Today is 2021-10-13.
Assistant: [AddAlarm(token="z9x8c7v6b5n4m3q2w1", time="2021-10-14 08:00:00")]
Tool: [AddAlarm Response: ”success"]
Assistant: An alarm has been set for 8 am tomorrow.
Figure 14: Example of API-Bank benchmark. The blue text represents the part to be evaluated.
System: You are an expert in composing functions. You are given a question and a set of possible functions. Based on the question, you will need to make one or more function/tool calls to achieve the purpose. If none of the function can be used, point it out. If the given question lacks the parameters required by the function, also point it out. You should only return the function call in tools call sections.
If you decide to invoke any of the function(s), you MUST put it in the format of [func_name1(params_name1=params_value1, params_name2=params_value2,...), func_name2(params)] You SHOULD NOT include any other text in the response. Here is a list of functions in JSON format that you can invoke. [{’name’: ’calc_absolute_pressure’, ’description’: ’Calculates the absolute pressure from gauge and atmospheric pressures.’, ’parameters’: { ’type’: ’dict’, ’properties’: { ’atm_pressure’: { ’type’: ’integer’, ’description’: ’The atmospheric pressure in atmospheres (atm). Default is 1 atm if not provided.’ }, ’gauge_pressure’: { ’type’: ’integer’, ’description’: ’The gauge pressure in atmospheres (atm). Must be provided.’ } }, ’required’: [ ’gauge_pressure’ ] }]
User: Calculate the absolute pressure in pascals given atmospheric pressure of 1 atm and a gauge pressure of 2 atm.
Assistant: [calc_absolute_pressure(atm_pressure=1, gauge_pressure=2)]
Figure 15: Case: model fine-tuned with our ToolACE data can solve the question in the zero-shot setting.
System: You are an expert in composing functions. You are given a question and a set of possible functions. Based on the question, you will need to make one or more function/tool calls to achieve the purpose. If none of the function can be used, point it out. If the given question lacks the parameters required by the function, also point it out. You should only return the function call in tools call sections.
If you decide to invoke any of the function(s), you MUST put it in the format of [func_name1(params_name1=params_value1, params_name2=params_value2,...), func_name2(params)] You SHOULD NOT include any other text in the response. Here is a list of functions in JSON format that you can invoke. [{"name": "solve_quadratic", "description": "Solve a quadratic equation given coefficients a, b, and c. If optional ’root_type’ is ’real’, the function will only return real roots. If not specified, function may return complex roots.", "parameters": {"type": "dict", "properties": {"a": {"type": "integer", "description": "The coefficient of the squared term in the quadratic equation."}, "b": {"type": "integer", "description": "The coefficient of the linear term in the quadratic equation."}, "c": {"type": "integer", "description": "The constant term in the quadratic equation."}, "root_type": {"type": "string", "description": "The type of roots to return: ’real’ for real roots, ’all’ for both real and complex roots. Default value is ’real’."}}, "required": ["a", "b", "c"]}}]
Here are some examples you can refer:
Available tools: [{’name’: ’FunctionIntersect.calculateRoots’, ’description’: ’Identifies the roots of the equation formed by setting two functions equal to each other.’, ’parameters’: {’type’: ’dict’, ’properties’: {’equation’: {’description’: "The equation obtained by setting two functions equal, e.g., ’3x^2 + 2x - 1 = x^3 - 2x + 4’.", ’type’: ’string’}, ’precision’: {’description’: ’The numerical precision for calculating roots, specified as a decimal.’, ’type’: ’float’}, ’method’: ’description’: "The numerical method to use for finding roots, such as ’Newton-Raphson’ or ’Bisection’.", ’type’: ’string’}}, ’required’: [’equation’]}}] User: Could you please help me find the roots of the equation 3x^2 - 7x + 2 = 0 using the Bisection method and again using the Newton-Raphson method? Assistant: [FunctionIntersect.calculateRoots(equation="3x^2 - 7x + 2 = 0", method="Bisection", precision=0.01), FunctionIntersect.calculateRoots(equation="3x^2 - 7x + 2 = 0", method="Newton-Raphson", precision=0.01)]
Available tools: [{’name’: ’complex_root_finder’, ’description’: ’Finds roots of a complex polynomial within a specified region.’, ...
User: I need to find the roots of the polynomial "z^4 + z^3 + 2z^2 + z + 6" within four different regions specified by the following centers and radii: (1+2i, 5), (-3+4i, 3), (0-1i, 4), (2+2i, 6). Each analysis should be done with high resolution during the optimal times of day.
Assistant: [complex_root_finder(polynomial="z^4 + z^3 + 2z^2 + z + 6", region={"center": {"real": 1.0, "imaginary": 2.0, "radius": 5.0}, search_parameters={"resolution": 1000, "time": "morning"}), complex_root_finder(polynomial="z^4 + z^3 + 2z^2 + z + 6", region={"center": {"real": -3.0, "imaginary": 4.0}, "radius": 3.0}, search_parameters={"resolution": 1000, "time": "morning"}), complex_root_finder(polynomial="z^4 + z^3 + 2z^2 + z + 6", region="center": "real": 0.0, "imaginary": -1.0, "radius": 4.0, search_parameters="resolution": 1000, "time": "morning"), complex_root_finder(polynomial="z^4 + z^3 + 2z^2 + z + 6", region={"center": {"real": 2.0, "imaginary": 2.0}, "radius": 6.0}, search_parameters={"resolution": 1000, "time": "morning"})]
[…]
User: I’m currently preparing for my calculus exam and need to solve some integrals. Can you solve the integral of the function f(x) = 3x^2 from x = 0 to x = 4 for me? Assistant: [calculus.integralSolver(function="lambda x: 3*x**2", limits={"lower": "0", "upper": "4"})]
$=$ User: Find all the roots of a quadratic equation given coefficients a = 3, b = -11, and c = -4.
Assistant: [calculus.integralSolver(function="lambda x: 3*x**2", limits="lower": "0", "upper": "4")]
Figure 16: Case: Few-shot in-context learning cannot solve the quite easy question, hallucinating on calling tools in few-shot examples instead of the test sample.